OpenStudy (anonymous):

(22x^5y^3z/16x^3z^5) times (8y^2z^3/11x^3)

OpenStudy (gw2011):

The first thing to do is to simplify this problem. (22x^5) (5y^3) (3z)/(16x^3) (3z^5) = (11x^2) (5y^3)/8z^4 Now the problem will look like this: [(11x^2) (5y^3)/8z^4] [(8y^2) (2z^3)/11x^3] The [ ] indicates that the knowns and unknows in the first [ ] is multiplied by the knowns and unknowns in the second [ ] Therefore, simplifying the above [ ] you get the following: [(5y^3)/z] [(2y^2)/x] = (5y^3) (2y^2)/(z) (x) = 10y^5/xz The answer is: 10y^5/xz